$SPAD/src/input mkfunc.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool mkfunc.output
)set message test on
)set message auto off
)clear all
--S 1 of 9
expr := (x - exp x + 1)**2 * (sin(x**2) * x + 1)**3
--R
--R
--R   (1)
--R       3   x 2        4     3   x    5     4    3      2 3
--R     (x (%e )  + (- 2x  - 2x )%e  + x  + 2x  + x )sin(x )
--R   +
--R        2   x 2        3     2   x     4     3     2      2 2
--R     (3x (%e )  + (- 6x  - 6x )%e  + 3x  + 6x  + 3x )sin(x )
--R   +
--R            x 2        2        x     3     2           2       x 2
--R     (3x (%e )  + (- 6x  - 6x)%e  + 3x  + 6x  + 3x)sin(x ) + (%e )
--R   +
--R                 x    2
--R     (- 2x - 2)%e  + x  + 2x + 1
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 9
function(expr, f, x)
--R
--R
--R   (2)  f
--R                                                                 Type: Symbol
--E 2
--S 3 of 9
tbl := [f(0.1 * i - 1) for i in 0..20];
--R
--R   Compiling function f with type Float -> Float
--R
--R                                                            Type: List(Float)
--E 3

--S 4 of 9
e := (x - y + 1)**2 * (x**2 * y + 1)**2
--R
--R
--R   (4)
--R      4 4        5     4     2  3     6     5    4     3     2      2
--R     x y  + (- 2x  - 2x  + 2x )y  + (x  + 2x  + x  - 4x  - 4x  + 1)y
--R   +
--R        4     3     2               2
--R     (2x  + 4x  + 2x  - 2x - 2)y + x  + 2x + 1
--R                                                    Type: Polynomial(Integer)
--E 4
--S 5 of 9
function(e, g, [x, y])
--R
--R
--R   (5)  g
--R                                                                 Type: Symbol
--E 5
--S 6 of 9
function(e, h, x, y)
--R
--R
--R   (6)  h
--R                                                                 Type: Symbol
--E 6
--S 7 of 9
m1 := squareMatrix [[1, 2], [3, 4]]
--R
--R
--R        +1  2+
--R   (7)  |    |
--R        +3  4+
--R                                                Type: SquareMatrix(2,Integer)
--E 7
--S 8 of 9
m2 := squareMatrix [[1, 0], [-1, 1]]
--R
--R
--R        + 1   0+
--R   (8)  |      |
--R        +- 1  1+
--R                                                Type: SquareMatrix(2,Integer)
--E 8
--S 9 of 9

h(m1, m2)
--R
--R   Compiling function h with type (SquareMatrix(2,Integer),SquareMatrix
--R      (2,Integer)) -> SquareMatrix(2,Integer)
--R
--R        +- 7836   8960 +
--R   (9)  |              |
--R        +- 17132  19588+
--R                                                Type: SquareMatrix(2,Integer)
--E 9
)spool
)lisp (bye)

———-

References^{[1] nothing}

6